# Information entropy for central 197Au+197Au collisions in the UrQMD model

### v1

@@ IV Conclusions
In summary, we simulated central 197Au+197Au collisions in the framework of UrQMD model which considers a standard mode or hybrid modes with three kinds of EoSs, i.e. EoS:HG without deconfinement transition, EoS:CH with the first-order phase transition and critical endpoint, and EoS:BM with a strong first-order phase transition between QGP and hadronic phase. By calculating the multiplicity information entropies of hadrons, anti-hadrons, baryons, anti-barons, net-baryons and net-protons, we explore the possible phase transition and EoS effects for the hydrodynamics. The dependence of information entropy on collision energy reflects the relationship between the width of multiplicity probability distribution of particles and collision energy. With the help of information entropies of baryons and net-baryons, it is found that both information entropies have valleys at around = 20 GeV with hydro hybrid modes, which indicates the onset of hydrodynamics behavior in the model. In addition, an enhancement emerges around = 30 GeV with EoS:CH and EoS:BM which is consistent with recent experimental observations from as well as intermittency analysis which give enhancement around =27 GeV, that may suggest the CEP. In one word, the present work suggests information entropy could be taken as an alternative observable on the QGP phase transition, and we expect that it can be checked from the experimental side.
@@ Acknowledgements.
Authors thank Dr. Chen Zhong for maintaining the high-quality performance of Fudan supercomputing platform for nuclear physics and the discussions with Dr. Kai-Jia Sun. This work received partial supports from the National Natural Science Foundation of China under Contract Nos. , , , ,and , the Strategic Priority Research Program of CAS under Grant No. XDB34000000, the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008,the STCSM under Grant No. 23590780100, and the Natural Science Foundation of Shanghai under Grant No. 23JC1400200.References
@@
- [8] F. P. Li, L. G. Pand, X. N. Wang, Nucl. Tech. (in Chinese) 46, 040014 (2023) .
@@
- [12] Z. Fodor and S. D. Katz, JHEP 03, 014 (2002).
- [13] P. de Forcrand and O. Philipsen, Nucl. Phys. B 673, 170 (2003).
- [14] H. T. Ding, S.T. Li, J.H. Liu, Nucl. Tech. (in Chinese) 46, 040008 (2023).
- [15] Z. R. Zhu, Y. Q. Zhao, D. F. Hou, Nucl. Tech. (in Chinese) 46, 040007 (2023).
@@
- [25] C. Sasaki, B. Friman, and K. Redlich, Phys. Rev. Lett. 99, 232301 (2007).
- [26] M. A. Stephanov, Phys. Rev. Lett. 102, 032301 (2009).
@@
- [35] K. J. Sun, L. W. Chen, C. M. Ko, et al., Nucl. Tech. (in Chinese) 46, 040012 (2023).
- [36] Dingwei Zhang (STAR Collaboration), Nucl. Phys. A 1005, 121825 (2021).
- [38] E. Shuryak and J. M. Torres-Rincon, Phys. Rev. C 101, 034914 (2020).
- [39] K. J. Sun and C. M. Ko, arXiv:2005.00182 [nucl-th] (2020).
- [40] X. G. Deng and Y. G. Ma, Phys. Lett. B 808, 135668 (2020).
@@
- [47] J. Steinheimer et al., Phys. Rev. C 77, 034901 (2008).
- [48] S. Gope and B. Bhattacharjee, Eur. Phys. J. A 57, 44 (2021).
- [49] H. C. Song, Y. Zhou, K. Gajdoov, Nucl. Sci. Tech. 28, 99 (2017).
- [50] Chun Shen, Li Yan, Nucl. Sci. Tech. 31, 112 (2020).
- [51] Z. Yang, T. Luo, W. Chen et al., Phys. Rev. Lett. 130, 052301 (2023).
- [52] Guang-You Qin, Nucl. Sci. Tech. 34, 22, (2023).
- [53] K. Yamazaki, T. Matsui, and G. Baym, Nucl. Phys. A 933, 245255, (2015).
- [54] C. Herold et al., Phys. Lett. B 790, 557–562 (2019).
@@
- [57] Chun-Wang Ma, Yu-Gang Ma, Prog. Part. Nucl. Phys. 99, 120 (2018).
- [58] F. Li and G. Chen, Eur. Phys. J. A 56:167 (2020).
- [59] M. Reiter et al, Nucl. Phys. A, 643, 29112, (1998).
- [60] A. Kovner, M. Lublinsky, and M. Serino, Phys. Lett. B 792, 4–15 (2019).
@@
- [78] Q. F. Li, J. Steinheimer, H. Petersen, et al., Phys. Lett. B 674, 111-116 (2009).
- [79] K. Morita, B. Friman, K. Redlich, Phys. Lett. B 714,178–183 (2015).

### v2

@@ IV Conclusions
In summary, we simulated central 197Au+197Au collisions using the UrQMD model considering both a standard mode and hybrid modes with three types of EoSs, i.e. EoS:HG without deconfinement transition, EoS:CH with the first-order phase transition and critical endpoint, and EoS:BM with a strong first-order phase transition between the QGP and hadronic phases. By calculating the multiplicity information entropies of hadrons, anti-hadrons, baryons, anti-baryons, net-baryons and net-protons, we investigate the possible phase transition and EoS effects for the hydrodynamics. The dependence of information entropy on collision energy illustrates the relationship between the width of the multiplicity probability distribution of particles and collision energy. With the help of baryon and net-baryon information entropies, it is observed that both exhibit valleys around GeV in the hydro hybrid modes, signaling the onset of hydrodynamic behavior in the model. Additionally, an enhancement emerges around = 30 GeV with EoS:CH and EoS:BM which is consistent with recent experimental observations from as well as intermittency analysis which give enhancement around =27 GeV, that may suggest the CEP. In one word, the present work suggests that the information entropy could serve as an alternative observable on the QGP phase transition.
@@ Acknowledgements.
Authors thank Dr. Chen Zhong for maintaining the high-quality performance of Fudan supercomputing platform for nuclear physics and the discussions with Dr. Kai-Jia Sun. This work received partial support from the National Natural Science Foundation of China under Contract Nos. , , , and ,, the Strategic Priority Research Program of CAS under Grant No. XDB34000000.References
- [1] P. Braun-Munzinger, V. Koch, T. Schäfer, J. Stachel, Phys. Rep. 621, 76 (2016).
- [2] W. Busza, K. Rajagopal, W. van der Schee, Ann. Rev. Nucl. Part. Sci. 68, 339 (2018).
- [3] A. Bzdak, S. Esumi, V. Koch, J. F. Liao, M. Stephanov, N. Xu, Phys. Rep. 853, 1 (2020).
- [4] A. Pandav, D. Mallick, B. Mohanty, Prog. Part. Nucl. Phys. 125, 103960 (2022).
- [5] J. H. Chen, X. Dong, X. He et al., Nucl. Sci. Tech. 25, 214 (2024).
- [6] Q. Y. Shou, Y. G. Ma, S. Zhang et al., arXiv:2409.17964.
@@
- [13] F. P. Li, L. G. Pang, X. N. Wang, Nucl. Tech. (in Chinese) 46, 040014 (2023) .
@@
- [26] J. Adam et al. (STAR Collaboration), Phys. Rev. Lett. 126, 092301 (2021).
- [27] C. Sasaki, B. Friman, and K. Redlich, Phys. Rev. Lett. 99, 232301 (2007). Phys. Rev. Lett. 102, 032301 (2009).
- [28] C. Liu, X. G. Deng, Y. G. Ma, Nucl. Sci. Tech. 33, 52 (2022).
@@
- [42] M. Bleicher, E. Bratkovskaya, Prog. Part. Nucl. Phys. 122, 103920 (2022).
@@
- [63] Q. F. Li, J. Steinheimer, H. Petersen, et al., Phys. Lett. B 674, 111 (2009).
- [65] Chun-Wang Ma, Yu-Gang Ma, Prog. Part. Nucl. Phys. 99, 120 (2018).
- [66] F. Li and G. Chen, Eur. Phys. J. A 56:167 (2020).
- [67] J. Pu, Y. B. Yu, K. X. Kai et al., Phys. Lett. B 841, 137909 (2023).
- [68] K. Kashiwa, H. Kouno, Phys. Rev. D 105, 054017 (2022).
- [69] W. Kou, X. Chen, Phys. Lett. B 846,138199 (2023).
- [70] H. L. Wei, X. Zhu, C. Yuan, Nucl. Sci. Tech. 33, 111 (2022).
- [71] C. Tsallis, J. Stat. Phys. 52, 479-487 (1988).
- [72] P. Jizba, T. Arimitsu, Annals of Physics 312, 17-59 (2004).
- [73] K. Morita, B. Friman, K. Redlich, Phys. Lett. B 714, 178 (2015).